OpenStudy (anonymous):

********** Metal ********** draw a triangle ABC with and angle C a right triangle, write sin A and cos B in terms of the side length of the triangle . what do you notice? how are angle A and angle B related? make a conjecture based on your observations??

OpenStudy (anonymous):



OpenStudy (anonymous):

sin (angle) = opposite side / hypotenuse sin A = BC / BA sin B = AC / BA Notice that they have similar factor of hypotenuse BA I also notice that A + B + 90 = 180, thus A+B =90

OpenStudy (anonymous):

@dauspex explain how the tangent of an acute angle changes as the angle measure increases please help :) @dauspex

OpenStudy (anonymous):

tan (angle) = number As angle increases, the tan value increases for acute (<90)

OpenStudy (anonymous):

thnx